OpenStudy (anonymous):

Solve the system by using elimination or elimination with multiplication. One point for the steps and one point for the correct solution x - y = 3 and x + y = 7

OpenStudy (mertsj):

add the two equations.

OpenStudy (mertsj):

What did you get?

OpenStudy (mertsj):

Ee. Let pdomingo do it

OpenStudy (anonymous):

This is a website where you ask other people to do you homework for you. I'm just trying to oblige.

OpenStudy (mertsj):

This is a website where you can get help when you need it.

OpenStudy (anonymous):

It was not intended to be that way.

OpenStudy (mertsj):

There are always people who will misuse everything and try to spoil it for everyone else.

OpenStudy (anonymous):

x - y = 3 x + y = 7 2x = 10 x = 5 y + 5 = 7 y = 2

OpenStudy (mertsj):

Congratulations. You have stolen from pdomingo the opportunity to do it for himself.

OpenStudy (anonymous):

Thanks man, I doubt hes even here anymore.

OpenStudy (anonymous):

he is actually her btw my name is pala domingo:) and thanks for the answer i just needed some help im not good in math at all and i didnt understand but now i do so yeah..

OpenStudy (anonymous):

No problem